# A new model of spontaneous scalarization of black holes induced by curvature and matter

### v1

@@ II.1 Description of the model
We consider a family of theories [^2] These theories can be considered as a generalization of Einstein-Maxwell-scalar-Gauss-Bonnet theories, and a subclass of Maxwell-Horndeski theories. where a real scalar field couples non-minimally both to the Gauss-Bonnet invariant and to a $U(1)$ gauge field (e.g. an electromagnetic field).
The action is :
@@
with $R$ the Ricci scalar and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ where $A_{\mu}$ is a $U(1)$ gauge field, $F_{\mu\nu}$ is its strength tensor (that can be identified as a Maxwell tensor), $\phi$ a scalar field non-minimally coupled to gravity and $U(1)$ gauge field through the functions $f(\phi)$ and $g(\phi)$ respectively, and $\mathcal{G}$ is the Gauss-Bonnet invariant which is defined as follows:
@@
  • Klein-Gordon equation:
    $\Box\phi+\frac{1}{4}f_{,\phi}(\phi)\mathcal{G}-\frac{1}{4}g_{,\phi}(\phi)F_{\mu\nu}F^{\mu\nu}=0$ (7)
@@ II.2 Field equations and ansatz
where $m(r)$ is the Misner-Sharp mass function [147] and $\sigma(r)$ a second metric function. Spherical symmetry, in the absence of a magnetic charge [^3] A magnetic charge would also be compatible with spherical symmetry, but it shall not be considered here - see [70] for dyonic charged BHs in the context of EMS theories., imposes an electrostatic 4-vector potential, $A=V(r)\,dt$, and a scalar field that depends only on $r$, $\phi=\phi(r)$.
Substituting into the field equations leads to a set of three coupled differential equations for the radial functions $N$, $\sigma$, and $\phi$, with an additional constraint. This system of equations can be put in a form suitable for numerical integration:
$N^{\prime}=F_{1}(N,\sigma,\phi,\phi^{\prime},\phi^{\prime\prime}),\qquad$ $\sigma^{\prime}=F_{2}(N,\sigma,\phi,\phi^{\prime},\phi^{\prime\prime}),$ (10)
$\phi^{\prime\prime}=F_{3}(N,\sigma,\phi,\phi^{\prime}),\quad$ $(\frac{V^{\prime}g(\phi)r^{2}}{\sigma})^{\prime}=0.$
@@ II.3 Boundary conditions and physical quantities of interest
To solve the set of four coupled ordinary differential equations (11), one must implement the appropriate boundary conditions. We impose regularity at the horizon, $r=r_{H}$, the field equations can be approximated by a power series expansion in $r-r_{H}$ as
@@
where the coefficients are given by
@@
The coefficients depend on the two parameters $\phi_{0}$ and $\sigma_{0}$, where the subscript 00 denotes functions evaluated at the horizon $r_{H}$.
@@
At spatial infinity we expand in power series of $1/r$, so the leading order expansion for far fields is
@@
where $\beta>0$ is an arbitrary scaling parameter, $\eta$ is the coupling constant of the GB term, defined so that $f=\eta\,\tilde{f}$, with dimensionless $\tilde{f}$, so that $\eta$ has dimension of length square.
This symmetry and freedom in choosing the length scale implies that physical quantities of the black hole solution should be expressed in terms of dimensionless combinations that remain invariant under these transformations. To characterize the black hole solutions we introduce the reduced quantities :
@@
II.4 Conditions for scalarization: Scalarization threshold and coupling functions
@@
Spontaneous scalarization occurs when the theory admits solutions with $\phi=0$. This demands
@@
A transition from a scalar-free black hole to one that supports a nontrivial scalar can occur if the second derivatives of $f(\phi)$ and $g(\phi)$ near $\phi=0$ create a tachyonic instability. Small-$\phi$ expansions for both coupling functions show:
@@
The linearized the Klein-Gordon equation then becomes
@@
and $\mu_{\rm eff,f}^{2}$ and $\mu_{\rm eff,g}^{2}$ are the effective mass squares.
A tachyonic instability appears when $\mu_{\rm eff}^{2}$ is negative, and:
@@
When $\mu_{\rm eff}^{2}<0$ a new branch of scalarized solutions bifurcate from the original scalar-free solutions (GR black holes). These scalarized solutions form a family of solutions which are continuously connected to the scalar-free solutions and characterized by additional parameters (e.g., the scalar field’s value at the horizon). The quantities $\mathcal{I}_{GB}$ and $\mathcal{I}_{M}$ influence both the onset of stability and the resulting scalarized solutions.
The choice of the coupling functions $f(\phi)$ and $g(\phi)$ plays an important role. Various forms have been proposed in the literature, which satisfy the condition that $f(0)=0$, $g(0)=1$, so that for vanishing scalar field one recovers the Einstein-Maxwell action, whose unique spherically symmetric solution is the Reissner-Nordström (RN) metric:
@@
$\mathcal{I}_{GB}=\frac{8}{r^{8}}\left[6M^{2}r^{2}-12Q^{2}Mr+5Q^{4}\right]$ (30)
@@
$\delta\phi=U(r)Y_{\ell m}(\theta,\varphi)$ (32)
This allows the scalar field perturbations equation (linearized KG equation) to be only radially dependent and written in the form
$\left[r^{2}N(r)U^{\prime}(r)\right]^{\prime}=U(r)(\ell(\ell+1)-\frac{\eta}{r^{2}}\left(\frac{12M^{2}}{r^{2}}+\frac{10Q^{4}}{r^{4}}-\frac{24MQ^{2}}{r^{3}}\right)+\alpha\frac{Q^{2}}{r^{2}}),$ (33)
@@
For $l=0$ this equation can be written as
@@
To facilitate our investigation, we introduce dimensionless variables :
@@
leading to the dimensionless form of the equation:
@@
The differential equation (36) can be solved numerically. We start our numerical integration near the event horizon (at $\rho=\rho_{H}+10^{-5}$), with $\rho_{H}=1+\sqrt{1-q^{2}}$  [^5] We will integrate the equation for the range of values $0\leqslant q\leqslant 1$; $q=1$ is the extremal RN black holes. using a power series solution for $u$ valid in that region. The leading-order behavior near the horizon is given by:
@@
We extend the integration to a large value of $\rho$, typically around $10^{5}$, for specific values of the coupling constants $(\gamma,\,\alpha)$ and an arbitrary horizon value $u_{H}$. For every set of these parameters, the equation (36) yields a distinct solution. Our focus is on solutions $u$ that vanish at spatial infinity.
@@
We varied these parameters systematically to map out the region in the parameter space where the GR is unstable in order to identify where scalarization can happen. This region is characterized by the onset of tachyonic instability, indicated by the conditions on the effective mass squared, $\mu_{\rm eff}^{2}$, as outlined in Eq. (23).
@@
In Fig. 1 we plot, for six fixed values of the matter coupling $\alpha$, the minimum dimensionless Gauss-Bonnet strength $\gamma=\eta/M^{2}$ that triggers the instability of a Reissner–Nordström black hole as a function of the charge ratio $q=Q/M$. The six panels are arranged by $\alpha$ from the top left $(\alpha=0)$ to the bottom right $(\alpha=-40)$. Each curve therefore marks the bifurcation line where a non-trivial scalar configuration first appears. The bullet list that follows comments on the behavior in every panel.
@@
  • Top-right, $\alpha=-2$ and $\alpha=-5$. The blue $(\alpha=-2)$ and orange $(\alpha=-5)$ curves bend below the $\gamma=0$ axis at high charge. The Maxwell coupling therefore allows scalarization with $\eta<0$ once $q$ is large.
  • Middle-left, $\alpha=-10$. The fundamental branch crosses $\gamma=0$ at $q\approx 0.6$, extending the $GB^{-}$ window that earlier work found only near $q\approx 0.96$ [85, 86]. Other branches appear, numerical checks confirm that the curve that appears just above the first fundamental branch is the first excited scalar mode with one radial node $(n=1)$.
[…]
@@
Overall, increasing $|\alpha|$ widens the parameter space in which the instability occurs and introduces additional node branches. It permits $\eta<0$ solutions with less restrictive conditions. The critical value $q^{(-)}$ at which $\eta$ flips sign decreases monotonically with $|\alpha|$; larger couplings push $\eta<0$ scalarization window down to $q\approx 0.2$, making negative-$\eta$ scalarization accessible at a much broader class of RN BH.
Earlier studies located this so-called $GB^{-}$ branch in narrow regimes. In case of spherical symmetry studied in [85, 86] RN clouds with $\eta<0$ were found only near extremality $q\approx 0.957058$. For rotating black holes, [148, 149, 150] reported the so-called spin-induced scalarization with $\eta<0$ scalarization when the spin is high ($J/M^{2}\gtrsim 0.5$. In both cases the Gauss-Bonnet invariant turns negative.
@@ III.1 Numerical procedure
We solve the system of ordinary differential equations in (11) with the boundary conditions in (12) and (14). These equations describe a boundary value problem with conditions at the event horizon and at large distance. Standard scaling arguments allow to set the horizon radius to $r_{H}=1$.
@@ III.2 Scalarized black hole solutions
In the following, we focus only on the case where $\eta$ is positive and $\alpha$ negative (which is the usual choice of sign for scalarization both in EsGB and EMS). One of the most important quantities to study for scalarized black holes is the scalar charge, which gives a clear idea of how scalarized black holes bifurcate from the GR ones and how they evolve.
@@
Figure 2 shows the plots of the scalar charge $\frac{D}{\sqrt{\eta}}$ with respect to the mass $\frac{M}{\sqrt{\eta}}$ for fixed ratio $Q/\sqrt{\eta}$, comparing several solutions with different $\alpha$. The scalar charge $\frac{D}{\sqrt{\eta}}$ versus mass $\frac{M}{\sqrt{\eta}}$ plots for scalarized black holes show that each choice of the coupling parameter $\alpha$ yields a distinct branch of solutions bifurcating from RN at different threshold mass. In single-coupling models (pure EsGB or EsGB plus a Ricci term) those mass thresholds are identical. In the present model, the threshold mass depends strongly on $\alpha$ and on $Q/\sqrt{\eta}$ This allows a shift of the bifurcation mass $\tilde{M}_{th}$ over a wide interval: the largest onset masses are up to three to four times the smallest ones. Furthermore after the bifurcation point, some scalarized branches extend toward higher masses while others extend toward lower masses. The direction of this extension appears to correlate with stability.
@@
  • Branch behavior at low charge: At low charge fraction $Q/\sqrt{\eta}=0.2$, the direction in which scalarized branches extend after bifurcation depends on $\alpha$ Coupling values $\alpha=0,-2,-5$ produce branches that extend toward higher mass values beyond the threshold (rightward on the plot). In contrast $\alpha=-20,-30,-40$ yield branches extending toward lower masses. The case $\alpha=-10$ shows a critical turning point where the branch has a curve that changes direction when analyzed individually in a small interval of $M/\sqrt{\eta}$ (see Fig. 3).
@@
  • Branch behavior at higher charge: For larger charge to coupling ratio (e.g. $Q/\sqrt{\eta}=1.6$ or above), all the scalarized solutions branches tend to extend toward smaller masses.
  • Stability implications: The direction in which a scalarized branch evolves is believed to indicate its stability. Branches that extended toward smaller masses (decreasing $M/\sqrt{\eta}$ as the scalar charge grows) are likely stable solutions, whereas those that extend toward larger masses beyond the bifurcation point may be unstable. This interpretation is in line with suggestions of [90, 91], which means that introducing the Matter coupling may have a stabilizing effect.
  • Convergent endpoints: Several of the scalar charge curves for different $\alpha$ converge toward similar end points in the plane. Despite their different onset masses and trajectories, many branches approach nearly the same final scalar charge (and associated mass) at their termination. This suggests that different coupling strengths $\alpha$ can lead to scalarized black holes with comparable limiting properties.

### v2

@@ II.1 Description of the model
We consider a family of theories [^2] These theories can be considered as a generalization of Einstein-Maxwell-scalar-Gauss-Bonnet theories (EMsGB), and a subclass of Maxwell-Horndeski theories. where a real scalar field couples non-minimally both to the Gauss-Bonnet invariant and to a $U(1)$ gauge field (e.g. an electromagnetic field). The action is
@@
with $R$ the Ricci scalar and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ where $A_{\mu}$ is a $U(1)$ gauge field, $F_{\mu\nu}$ is its strength tensor (that can be identified as a Maxwell tensor), $\phi$ is a scalar field non-minimally coupled to gravity and the $U(1)$ gauge field through the functions $f(\phi)$ and $g(\phi)$ respectively, and $\mathcal{G}$ is the Gauss-Bonnet invariant which is defined as follows:
@@
  • Klein-Gordon (KG) equation:
    $\Box\phi+\frac{1}{4}f_{,\phi}(\phi)\mathcal{G}-\frac{1}{4}g_{,\phi}(\phi)F_{\mu\nu}F^{\mu\nu}=0\$(7)
@@ II.2 Field equations and ansatz
where $m(r)$ is the Misner-Sharp mass function [151] and $\sigma(r)$ is a second metric function. Spherical symmetry, in the absence of a magnetic charge [^3] A magnetic charge would also be compatible with spherical symmetry, but it shall not be considered here - see [70] for dyonic charged BHs in the context of EMS theories., imposes an electrostatic 4-vector potential, A=V(r)dtA=V(r)\,dt, and a scalar field that depends only on $r$, $\phi=\phi(r)$.
Substituting this ansatz into the field equations leads to a set of three coupled differential equations for the radial functions $N$, $\sigma$, and $\phi$, with an additional constraint. This system of equations can be put in a form suitable for numerical integration as
$N^{\prime}=F_{1}(N,\sigma,\phi,\phi^{\prime},\phi^{\prime\prime}),\qquad$ $\sigma^{\prime}=F_{2}(N,\sigma,\phi,\phi^{\prime},\phi^{\prime\prime}),$ $\phi^{\prime\prime}=F_{3}(N,\sigma,\phi,\phi^{\prime})$ (10)
$\qquad\text{and}\quad(\frac{V^{\prime}g(\phi)r^{2}}{\sigma})^{\prime}=0,$
@@ II.3 Boundary conditions and physical quantities of interest
To solve the set of four coupled ordinary differential equations Eq (11), one must implement the appropriate boundary conditions. We impose regularity at the horizon, $r=r_{H}$, and the field equations can be approximated by a power series expansion in $r-r_{H}$ as
@@
where the coefficients are obtained by using the expansion into the field equations:
@@
The coefficients depend on the two parameters $\phi_{0}$ and $\sigma_{0}$, where the subscript 0 denotes functions evaluated at the horizon $r_{H}$.
@@
At spatial infinity we expand in power series of $1/r$, so the leading-order expansion for far fields is
@@ II.4 Scaling symmetry
where $\beta>0$ is an arbitrary scaling parameter, and $\eta$ is the coupling constant of the GB term, defined so that $f=\eta\,\tilde{f}$, with dimensionless $\tilde{f}$, so that $\eta$ has the dimension of length square.
This symmetry and freedom in choosing the length scale implies that physical quantities of the black hole solution should be expressed in terms of dimensionless combinations that remain invariant under these transformations. To characterize the black hole solutions we introduce the reduced quantities
@@
II.5 Conditions for scalarization: Scalarization threshold and coupling functions
@@
Two conditions are required for spontaneous scalarization to happen, as follows:
  • When the theory admits solutions with $\phi=0$. This demands
@@
  • If the second derivatives of $f(\phi)$ and $g(\phi)$ near $\phi=0$ create a tachyonic instability. Small-$\phi$ expansions for both coupling functions show
@@
    The linearized Klein-Gordon equation then becomes
@@
    and $\delta\phi$ is the scalar field perturbation, while $\mu_{\rm eff,f}^{2}$ and $\mu_{\rm eff,g}^{2}$ are the effective mass squares.
    A tachyonic instability appears when $\mu_{\rm eff}^{2}$ is negative, and
@@
When $\mu_{\rm eff}^{2}<0$ a new branch of scalarized solutions may bifurcate from the original scalar-free solutions (GR black holes). These scalarized solutions form a family of solutions which are continuously connected to the scalar-free solutions and characterized by additional parameters (e.g., the scalar field’s value at the horizon). The quantities $\mathcal{I}_{GB}$ and $\mathcal{I}_{M}$ influence both the onset of stability and the resulting scalarized solutions.
The choice of the coupling functions $f(\phi)$ and $g(\phi)$ plays an important role. Various forms have been proposed in the literature, which satisfy the condition that $f(0)=0$, $g(0)=1$, so that for the vanishing scalar field one recovers the Einstein-Maxwell action, whose unique spherically symmetric solution is the Reissner-Nordström (RN) metric.
@@
$\mathcal{I}_{GB}=\frac{8}{r^{8}}\left[6M^{2}r^{2}-12Q^{2}Mr+5Q^{4}\right]\$(30)
@@
$\delta\phi=U(r)Y_{\ell m}(\theta,\varphi)\$(32)
This allows the scalar field perturbation equation (linearized KG equation) to be only radially dependent and written in the form
$\left[r^{2}N(r)U^{\prime}(r)\right]^{\prime}=U(r)\left(\ell(\ell+1)-\frac{\eta}{r^{2}}(\frac{12M^{2}}{r^{2}}+\frac{10Q^{4}}{r^{4}}-\frac{24MQ^{2}}{r^{3}})+\alpha\frac{Q^{2}}{r^{2}}\right).$ (33)
@@
For $\ell=0$ this equation can be written as
@@
To facilitate our investigation, we introduce dimensionless variables
@@
leading to the dimensionless form of the equation
@@
The differential equation Eq (36) can be solved numerically. We start our numerical integration near the event horizon (at $\rho=\rho_{H}+10^{-5}$), with $\rho_{H}=1+\sqrt{1-q^{2}}$  [^5] We will integrate the equation for the range of values $0\leqslant q\leqslant 1$, where $q=1$ is the extremal RN black hole. using a power series solution for $u$ valid in that region. The leading-order behavior near the horizon is given by:
@@
We extend the integration to a large value of $\rho$, typically around $10^{5}$, for specific values of the coupling constants $(\gamma,\,\alpha)$ and an arbitrary horizon value $u_{H}$. For every set of these parameters, Eq (36) yields a distinct solution. Our focus is on solutions $u$ that vanish at spatial infinity.
@@
We varied these parameters systematically to map out the region in the parameter space where the GR is unstable in order to identify where scalarization can happen. This region is characterized by the onset of tachyonic instability, indicated by the contribution of effective mass squared, $\mu_{\rm eff}^{2}$, as outlined in Eq. (33).
@@
In Fig. 1 we plot, for six fixed values of the matter coupling $\alpha$, the minimum dimensionless Gauss-Bonnet strength $\gamma=\eta/M^{2}$ that triggers the instability of a Reissner–Nordström black hole as a function of the charge ratio $q=Q/M$. The six panels are arranged by $\alpha$ from the top left $(\alpha=0)$ to the bottom right $(\alpha=-40)$. Each curve, therefore, marks the bifurcation line where a non-trivial scalar configuration first appears. The bullet list that follows comments on the behavior in every panel:
@@
  • Top-right, $\alpha=-2$ and $\alpha=-5$. The blue $(\alpha=-2)$ and orange $(\alpha=-5)$ curves bend below the $\gamma=0$ axis at high charge. The Maxwell coupling, therefore, allows scalarization with $\eta<0$ once $q$ is large.
  • Middle-left, $\alpha=-10$. The fundamental branch crosses $\gamma=0$ at $q\approx 0.6$, extending the $GB^{-}$ window that earlier work found only near $q\approx 0.96$ [85, 86]. Other branches appear, and numerical checks confirm that the curve that appears just above the first fundamental branch is the first excited scalar mode with one radial node $(n=1)$.
  • Middle-right, $\alpha=-20$. The sign flip now occurs near $q\approx 0.4$. Three excited branches (n=1, 2, 3)(n=1,\;2,\;3) are visible. Larger $|\alpha|$, therefore, lowers the charge threshold and creates more bifurcation lines.
@@
Overall, increasing $|\alpha|$ widens the parameter space in which the instability occurs and introduces additional node branches. It permits $\eta<0$ solutions with less restrictive conditions. The critical value $q^{(-)}$ at which $\eta$ flips sign decreases monotonically with $|\alpha|$, and larger couplings push the $\eta<0$ scalarization window down to $q\approx 0.2$, making negative-$\eta$ scalarization accessible at a much broader class of RN BH.
Earlier studies located this so-called $GB^{-}$ branch in narrow regimes. In case of spherical symmetry studied in [85, 86] RN clouds with $\eta<0$ were found only near extremality $q\approx 0.957058$. For rotating black holes, Refs [152, 153, 154] reported the so-called spin-induced scalarization with $\eta<0$ scalarization when the spin is high ($J/M^{2}\gtrsim 0.5$. In both cases the Gauss-Bonnet invariant turns negative.
@@ III.1 Numerical procedure
We solve the system of ordinary differential equations in (11) with the boundary conditions in (12) and (14). These equations describe a boundary value problem with conditions at the event horizon and at large distance. Standard scaling arguments allow us to set the horizon radius to $r_{H}=1$.
@@ III.2 Scalarized black hole solutions
In the following, we focus only on the case where $\eta$ is positive and $\alpha$ is negative (which is the usual choice of sign for scalarization both in EsGB and EMS). One of the most important quantities to study for scalarized black holes is the scalar charge, which gives a clear idea of how scalarized black holes bifurcate from the GR ones and how they evolve.
@@
Figure 2 shows the plots of the scalar charge $\frac{D}{\sqrt{\eta}}$ with respect to the mass $\frac{M}{\sqrt{\eta}}$ for fixed ratio $Q/\sqrt{\eta}$, comparing several solutions with different $\alpha$. The scalar charge $\frac{D}{\sqrt{\eta}}$ versus mass $\frac{M}{\sqrt{\eta}}$ plots for scalarized black holes show that each choice of the coupling parameter $\alpha$ yields a distinct branch of solutions bifurcating from RN at different threshold masses. In single-coupling models (pure EsGB or EsGB plus a Ricci term) those mass thresholds are identical. In the present model, the threshold mass depends strongly on $\alpha$ and on $Q/\sqrt{\eta}$ . This allows a shift of the bifurcation mass $\tilde{M}_{th}$ over a wide interval: the largest onset masses are up to three to four times the smallest ones. Furthermore after the bifurcation point, some scalarized branches extend toward higher masses while others extend toward lower masses. The direction of this extension appears to correlate with stability, as outlined below:
@@
  • Branch behavior at low charge: at low charge fraction $Q/\sqrt{\eta}=0.2$, the direction in which scalarized branches extend after bifurcation depends on $\alpha$ . Coupling values $\alpha=0,-2,-5$ produce branches that extend toward higher mass values beyond the threshold (rightward on the plot). In contrast $\alpha=-20,-30,-40$ yield branches extending toward lower masses. The case $\alpha=-10$ shows a critical turning point where the branch has a curve that changes direction when analyzed individually in a small interval of $M/\sqrt{\eta}$ (see Fig. 3).
@@
  • Branch behavior at higher charge: for larger charge to coupling ratio (e.g. $Q/\sqrt{\eta}=1.6$ or above), all the scalarized solution branches tend to extend toward smaller masses.
  • Stability implications: the direction in which a scalarized branch evolves is believed to indicate its stability. Branches that extended toward smaller masses (decreasing $M/\sqrt{\eta}$ as the scalar charge grows) are likely stable solutions, whereas those that extend toward larger masses beyond the bifurcation point may be unstable. This interpretation is in line with the suggestions of [90, 91], which means that introducing the matter coupling may have a stabilizing effect.
  • Convergent end points: several of the scalar charge curves for different values of $\alpha$ converge toward similar end points in the plane. Despite their different onset masses and trajectories, many branches approach nearly the same final scalar charge (and associated mass) at their termination. This suggests that different coupling strengths $\alpha$ can lead to scalarized black holes with comparable limiting properties. The behavior of the curves suggests maybe the existence of an attractor.